Can someone please help me make a systems of equations using this problem:
Your local cable television company offers two plans: basic service with one movie channel for $35 per month, or basic service with two movie channels for $45 per month. What is the charge for the basic service and the charge for each movie channel?

Found 2 solutions by FrankM, addingup:
Answer by FrankM(1040)

If it's linear (and few things in life are) we have
B+2M = 45
B+1M = 35
So, M =$10 and B = $25

Answer by addingup(3677)

movie channel: x
35+x = 45
x = 45-35
x = 10 this is the cost for a movie channel.